Circular No. 6000

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1993ak IN UGC 8685 H. Meusinger, Thüringer Landessternwarte, Tautenburg, reports the

discovery of an apparent supernova on a Tautenburg Schmidt plate taken by F. Börngen on 1993 Mar. 30.1 UT. The candidate was then at B = 16.5 and was located 25′′ east of the center of UGC 8685 = MCG 05-32-074 (α = 13^{h}40^{m}52^{s}.2, δ = +30^{o}35′16′′, equinox 1950.0). On a second Tautenburg plate taken 1993 Apr. 27.0, SN 1993ak had dimmed to B = 18.5. There is no stellar object at this position on several inspected plates taken before 1977 and after 1994 May 2.9. Plates taken under good seeing conditions show a relatively bright diffuse patch in a spiral arm (presumably a giant H II region) at the position of SN 1993ak.

GRS 1915+105 P. Durouchoux, S. Corbel, and P. Wallyn, Centre d’Etudes de Saclay,

Service d’Astrophysique; and J. Grindlay, Center for Astrophysics (CfA), communicate: “We observed two molecular clouds in the direction of the hard x-ray transient and possible soft γ-ray repeater (SGR) GRS 1915+105 on May 8, via the Swedish-European Southern Observatory Submillimeter Telescope at La Silla. Using transitions of ^{12}CO(1–0) at 115 GHz and 13CO(1–0) at 110 GHz, 10×10 grid maps were done centered on the radio counterpart (discovered by Mirabel et al. 1994, A.Ap. 282, 17) with an angular resolution of 45′′, a grid spacing of 45′′, and 90 s of integration time per position. We deduced, from a preliminary analysis, the following velocities, widths, and implied mean effective column densities: first cloud, v_{LSR} = 26 km/s, FWHM ∼ 5 km/s, N(H_{2}) ∼ 3 × 10^{21} cm^{−2}; second cloud, 59 km/s, 9 km/s, 2 × 10^{22} cm^{−2}. Both velocities are in good agreement with the previous Columbia CO survey (Dame et al. 1986, Ap.J. 305, 892) and also recent observations from the CfA (Grindlay 1994, Ap.J. Suppl., in press). This second cloud, with a measured kinematic distance of ∼ 7.8 kpc, exhibiting a larger line width and a greater column density (which is in good agreement with the column density deduced from ROSAT observations by Greiner et al. 1994, AIP Conf. Proc. No. 304, 260), is a good candidate for a possible association with GRS 1915+105. These results are consistent with the model proposed by Grindlay (loc. cit.), where SGR sources are due to thermal instabilities in spherical accretion onto black holes in giant molecular clouds. Higher-resolution maps (both radio and infrared), as well as spectra (to search for x-ray ionization effects), are needed.”

1994 June 3 Daniel W. E. Green